Find 7 ÷ 3.5 Find 7 ÷ 3.5 Find 7 ÷ 3.5 Find 7 ÷ 3.5

Find
 7 ÷ 3.5
 Find
 7 ÷ 3.5 
Find
 7 ÷ 3.5 
Find
 7 ÷ 3.5 

Grade:12

1 Answers

Harshit Singh
2 years ago
Dear Student

We have,
= 7 ÷ (35/10)
= 7 × (10/35)
= 1 × (10/5)
=2

Thanks